Grade 11Trigonometry

Prove that if one of the line of ax2+2hxy+by2=0 bisects the angle between axes then a+b=+-2h

Rowel
7 Years agoGrade 11

1 Answer

Rajat
7 Years ago
If one of the line bisects the angle between the perpendicular axes then it's eqn must be; x=y or x=(-y)  thus putting the value of X in the given eqn we get; ay^2±2hy^2+by^2=0
or, a ± 2h +b= 0
Or, a + b = ±2h 
Hence proved